Should the definition of random vector be associated with P: the probability set function?


1
Expert's answer
2021-09-27T16:42:13-0400

The concept of random vectors is a multidimensional generalization of the concept of random variable.

Suppose that we conduct a probabilistic experiment and that the possible outcomes of the experiment are described by sample space "\Omega" .

A random vector is a vector whose value depends on the outcome of the experiment, as stated by the following definition.

Let "\Omega" be a sample space. A random vector X is a function from sample space "\Omega" to the set of K-dimensional real vectors "\mathbb{R}^{\kappa}" :

"X = \Omega → \mathbb{R}^{\kappa}"

In rigorous probability theory, the function X is also required to be measurable. We report here a more rigorous definition of random vector by using the formulation of measure theory.

Let "(\Omega, F, P)" be a probability space. Let "\beta (\mathbb{R}^{\kappa})" be the Borel sigma- algebra of "\mathbb{R}^{\kappa }" . A function "X : \Omega \rightarrow \mathbb{R}^{\kappa }" such that "\left \{ \infty \epsilon \Omega :X(\infty ) \epsilon B\right \} \epsilon F \; for \; any \; B \epsilon \beta (\mathbb{R}^{\kappa })" is said to be a random variable on "\Omega" .

This definition ensures that the probability that realization of the random vector X will belong to set "B \epsilon \beta (\mathbb{R}^{\kappa})" can be defined as "P( X \epsilon B) = P({ \infty \epsilon \Omega : X(\infty ) \epsilon B})" because the set "{ \infty \epsilon \Omega : X(\infty ) \epsilon B}" belongs to the sigma- algebra F and as a consequence it's probability is well defined.